OpenStudy (anonymous):

(1+ i square root 3)^8? in standard form?

OpenStudy (anonymous):

i think the answer is 82..i like did this in my head

OpenStudy (anonymous):

(1+i√3)^8=(((1+i√3)^2)^2)^2

OpenStudy (anonymous):

because i^8 is 1 and radical(square root) 3 to the 8th is 81 so 81 plus 1 is 82

OpenStudy (anonymous):

Continuation of my previous post: (1+i√3)^8 (((1+i√3)^2)^2)^2 (1+i√3)^2=1+2i√3-3=2i√3-2=2(i√3-1) (2(i√3-1))^2=4(-3-2i√3+1)=4(-2-2i√3)=-8(1+i√3) (-8(1+i√3))^2=64(1+2i√3-3)=64(2i√3-2)=128(i√3-1) In standard form: -128+128i√3

OpenStudy (perl):

i would change first to trig form, then use demoivre's theorem

OpenStudy (perl):

if we plot the number inside the expresion, ( 1 + i sqrt(3)) , we get ( 1 , sqrt(3)) change this to trig form. r = sqrt( 1 ^2 + sqrt(3)^2) = sqrt( 1 + 3) + sqrt( 4) = 2 and tan(theta) = sqrt(3)/1 , theta = arctan sqrt(3) theta = 60 degrees

OpenStudy (perl):